OpenStudy (anonymous):

3x+y = 6.5 2x = 4y - 5

OpenStudy (♪chibiterasu):

Your question?

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (anonymous):

first step is: 3x - y = 6.5 4y - 2x = -5

OpenStudy (anonymous):

is that right?

OpenStudy (cggurumanjunath):

good

OpenStudy (cggurumanjunath):

what is next step ?

OpenStudy (anikate):

3x+y = 6.5 2x = 4y - 5 then, 3x+y=6.5 2x-4y=-5 then, (3x+y=6.5)2 (2x-4y=-5)3 then, 6x+2y=13 6x-12y=-15 then subtract, 14y=28 the, divide both sides by 14, 14y/14=28/14, so then the 14s would cross out and 28/14 is 2. so. y=2

OpenStudy (anikate):

then substitute y into an original expression in your question.

OpenStudy (cggurumanjunath):

3x - y = 6.5 - 2x+4y = -5 is ur next step right ?

OpenStudy (anonymous):

i dont know step 2 :/

OpenStudy (cggurumanjunath):

3x - y = 6.5 - 2x+4y = -5 is ur next step right ?====step 2

OpenStudy (anikate):

4(2) - 2x =-5 then, 8 - 2x =-5 then, -2x= -5 -8, if hope u know what i did there. -2x=-13, then dive both side by -2 x=13/2

OpenStudy (anonymous):

3x - y = 6.5 - 2x+4y = -5 this is step 2?

OpenStudy (cggurumanjunath):

yes

OpenStudy (cggurumanjunath):

what next ?

OpenStudy (anonymous):

(3x-y=6.5) - (4y-2x=-5)

OpenStudy (cggurumanjunath):

3x - y = 6.5====multiply this equation by the coeffient of x in the second equation ie -2 - 2x+4y = -5====multiply this equation by the coeffient of x in the first equation ie 3

OpenStudy (cggurumanjunath):

no this step will come in the end...(3x-y=6.5) - (4y-2x=-5)

OpenStudy (cggurumanjunath):

3x - y = 6.5====multiply this equation by the coeffient of x in the second equation ie -2 - 2x+4y = -5====multiply this equation by the coeffient of x in the first equation ie 3

OpenStudy (anonymous):

omg this is so hard i dont understand anything >,<

OpenStudy (cggurumanjunath):

-6x+2y=-13 -6x+12y=-15

OpenStudy (cggurumanjunath):

do one thing take 5 minutes and write down steps by steps the problem which i explained before ,then u can call me. ok////

OpenStudy (anonymous):

i wrote everything down already :/

OpenStudy (anonymous):

now whats after this? -6x+2y=-13 -6x+12y=-15

OpenStudy (cggurumanjunath):

-6x+2y=-13===call this equation 3 -6x+12y=-15===call this equation 4. do equation(3)-equation(4) now .

OpenStudy (cggurumanjunath):

tell me the answer @charlieberzak

OpenStudy (cggurumanjunath):

it is lengthy problem but it is easy.

OpenStudy (cggurumanjunath):

(-6x+2y=-13)-(-6x+12y=-15)

OpenStudy (anonymous):

(-6x+2y=-13)-(-6x+12y=-15) -6x--6x+2y+12y=-13-15

OpenStudy (anonymous):

-6x+6x+2y+12y = -13-15

OpenStudy (cggurumanjunath):

-6x+6x+2y+12y=-13+15

OpenStudy (cggurumanjunath):

be careful with the + and - signs.

OpenStudy (anonymous):

yes

OpenStudy (cggurumanjunath):

what next ?

OpenStudy (anonymous):

14xy = 28

OpenStudy (cggurumanjunath):

?

OpenStudy (cggurumanjunath):

14y=2 right ?

OpenStudy (anonymous):

i dont know :/

OpenStudy (cggurumanjunath):

-6x+6x+2y+12y=-13+15 2y+12y=2 right ?

OpenStudy (anonymous):

ummm not 14y?

OpenStudy (anonymous):

2y+12y=14y

OpenStudy (cggurumanjunath):

why ?

OpenStudy (cggurumanjunath):

it is 14y.

OpenStudy (anonymous):

because u said this -6x+6x+2y+12y=-13+15 2y+12y=2 right ?

OpenStudy (cggurumanjunath):

-13+15=2

OpenStudy (anonymous):

ohhh isee what u mean, never mind :)

OpenStudy (cggurumanjunath):

-13+15 is not 28

OpenStudy (cggurumanjunath):

be careful with the + and - signs.

OpenStudy (cggurumanjunath):

14y=2

OpenStudy (anonymous):

yes, iknow :)

OpenStudy (cggurumanjunath):

now find y.

OpenStudy (anonymous):

14y/-14 = 2/14

OpenStudy (anonymous):

is that how i should do it?

OpenStudy (cggurumanjunath):

14y/14=2/14

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

y=7

OpenStudy (cggurumanjunath):

why negative sign ? 14y/-14 = 2/14

OpenStudy (cggurumanjunath):

no

OpenStudy (cggurumanjunath):

y=2/14=1/7

OpenStudy (cggurumanjunath):

can u tell me x value now correctly ?

OpenStudy (cggurumanjunath):

y=1/7 is it clear ?

OpenStudy (anonymous):

7=1/7

OpenStudy (cggurumanjunath):

7=1/7 how ?

OpenStudy (cggurumanjunath):

n0 7=1/7 is wrong.

OpenStudy (anonymous):

omfg i dont know :/ i suck at this ****

OpenStudy (cggurumanjunath):

14y/14=2/14

OpenStudy (cggurumanjunath):

do u agree with above statement ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

14y/14 = y

OpenStudy (anonymous):

correct?

OpenStudy (cggurumanjunath):

yes

OpenStudy (cggurumanjunath):

2/14

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

7

OpenStudy (cggurumanjunath):